OpenStudy (anonymous):

y^5/x/y^2/x^2 in simplest form? I got x/y^3... Is that right??

OpenStudy (anonymous):

Anyone understand this?

OpenStudy (aroub):

y^3/x^3 u mean?

OpenStudy (anonymous):

Wait for the question or answer?

OpenStudy (phi):

Is this what you mean? \[\frac{\frac{y^{5}}{x}}{\frac{y^{2}}{x^{2}}}\]

OpenStudy (anonymous):

Yes :)

OpenStudy (aroub):

answer...

OpenStudy (phi):

Multiply the top and bottom fraction by x^2/y^2 The bottom fraction turns into 1.

OpenStudy (phi):

The top is \[\frac{y^{5}}{x}\cdot \frac{x^{2}}{y^{2}}\] Can you finish the problem?

OpenStudy (anonymous):

Thank you :)